OpenStudy (anonymous):

I had a question about using the U substitution for finding antiderivatives:

OpenStudy (anonymous):

\[\int\limits_{}^{?}(e^x+e ^{-x})/2\]

OpenStudy (anonymous):

That 2 is supposed to be inside the parenthesis, but I didn't know how to do that. So anyways, would a good approach to this be using the e^-x as your U?

zepdrix (zepdrix):

What do you mean by... inside the parenthesis? :O I don't think that really changes the problem :D

OpenStudy (anonymous):

Oh! Really? I didn't know, sorry. I thought that would've changed the problem entirely.

zepdrix (zepdrix):

\[\large \int\limits \frac{e^x+e^{-x}}{2}dx\]Pull the constant outside, then write it as 2 separate integrals, maybe that will make it easier to work with :)\[\large \frac{1}{2}\int\limits e^x+e^{-x}dx\qquad \rightarrow \qquad \frac{1}{2}\int\limits e^x dx+\frac{1}{2}\int\limits e^{-x} dx\]Now you should only really worry about doing a U sub for the second integral. The first one is pretty straight forward :D

OpenStudy (anonymous):

Right, okay lemme see if I get this.

OpenStudy (anonymous):

Oh wait...okay so basically you don't need to use the U sub. then, correct?

zepdrix (zepdrix):

If you are familiar with taking the derivative of an exponential, then you'll very quickly get familiar with the anti-derivative of an exponential. Unless you already understand the process? :) You certainly don't need a U-sub, unless you still need to learn this concept with what happens to the constant on the exponent. the negative one in this case.

OpenStudy (anonymous):

Can you show me that concept, please? I think I may have forgotten. Is the answer for that just -e^-x?

zepdrix (zepdrix):

Yes. Since it's just a negative, it's pretty straight forward. But lemme give a quick example with a different constant so it will sink in. :D

OpenStudy (anonymous):

Thank you so much, I appreciate this.

zepdrix (zepdrix):

Sorry it takes me a little bit to type stuff :C the equation tool CONSTANTLY crashes my browser...

zepdrix (zepdrix):

\[\large \int\limits e^{2x}dx\]Hmmm let's say we don't know this integral. Let's start by taking the derivative of this term and see what happens.\[\huge (e^{2x})'\quad =\quad e^{2x}(2x)'\quad=\quad2e^{2x}\]Hmm ok. Let's take the integral of the thing we just differentiated. \[\huge \int\limits\limits 2e^{2x}dx \quad = \quad e^{2x}\]How did we end up solving that integral? We actually DIVIDED by 2 didn't we? See how the 2 disappeared?

OpenStudy (anonymous):

Oooh! Okay that makes sense.

zepdrix (zepdrix):

So we divided by 2... now let's return to the integral we started with.\[\huge \int\limits e^{2x}dx\]The fact that there isn't a 2 in front shouldn't affect the integration process at all. So that's basically the rule for exponentials, you DIVIDE by the coefficient on the exponent.

zepdrix (zepdrix):

If the exponent is MORE than just x, let's say x^2, then that wouldn't apply :O If it's a constant attached to X though, then we're ok :D

zepdrix (zepdrix):

\[\huge \int\limits\limits e^{2x}dx=\frac{1}{2}e^{2x}\]Just in case I didn't make it clear :D heh

OpenStudy (anonymous):

Ah, okay. Thank you, I think I understand that now. So, what if it was like...say e^ln2? Would that apply? Sorry, I'm just throwing a random question in case I see it on my exam.

zepdrix (zepdrix):

Well try to be careful, sometimes you'll get things thrown at you that LOOK confusing, but are really just CONSTANTS. I like to call them fancy constants. They have their tuxedo on, they're really for the ball. But it's just an impostor, posing as something else. Example:\[\large (e^{\ln2})'=0\]See how it contains no X's? When we integrate, it will be a tiny bit different.

zepdrix (zepdrix):

It's similar to taking the integral of .... 3. An x pops back into the problem right?\[\large \int\limits 3 dx=3x\]\[\large \int\limits e^{\ln2}dx=e^{\ln2}x\]

OpenStudy (anonymous):

Oooh!!! Ha! That's really tricky, I guess that ln just messes me up sometimes. Wow :D

zepdrix (zepdrix):

If we had say,\[\large \int\limits e^{\ln2 x}dx\]Then yes, this rule that we went over earlier would still apply :)\[\frac{1}{\ln 2}e^{\ln 2 x}\]

zepdrix (zepdrix):

woops, the x is suppose to be outside of the natural log, i didn't write that very clearly.

OpenStudy (anonymous):

Okay so, basically, for that problem, you do the exact same thing as the other log problem so, you just divide that with the ln2? Right? I Think I'm getting it.

zepdrix (zepdrix):

Yah, when you DIFFERENTIATE, you end up MULTIPLYING by the coefficient on x. When you INTEGRATE, you DIVIDE by the coefficient on x.

zepdrix (zepdrix):

Here's a tricky one to try on for size :D\[\huge \int\limits e^{3x+1}dx\]

OpenStudy (anonymous):

Okay okay lemme see.

OpenStudy (anonymous):

...okay this is what I got...soooo would it be \[1\div3x+1\times e ^{3x+1}\] ?

zepdrix (zepdrix):

So the issue with this problem is that I threw that nasty +1 into the exponent! :O

zepdrix (zepdrix):

we gotta deal with that before we can integrate XD I so sneaky! lol

OpenStudy (anonymous):

Haha! WEll that's good! I have to learn this anyways :P

zepdrix (zepdrix):

Remember your rules of exponents?\[\huge e^{a+b}=e^a\cdot e^b\]Does this look familiar? :D

OpenStudy (anonymous):

Is it one of those log rules? It does look familiar

OpenStudy (anonymous):

Ooh! Or those rules when working with exponents.

zepdrix (zepdrix):

It's an exponent rule, but yah you prolly learned them at the same time you learned log rules :D

OpenStudy (anonymous):

Ooh! So...I'm guessing I had to split the problem first..and then integrate it?

zepdrix (zepdrix):

Yah :D

OpenStudy (anonymous):

Gah, that changes everything, okay okay lemme see if I got this now.

OpenStudy (anonymous):

Okay so I got \[\int\limits_{?}^{?} 1\div3x \times \int\limits_{?}^{?}e^1\]

OpenStudy (anonymous):

Then...it should just be 1/3x (e^3x) and then times that with e^1 which is....zero? Sorry I'm a bit shaky with those logs.

OpenStudy (anonymous):

Oh wait nope! It's 1!

OpenStudy (anonymous):

Oooh!!! no no, my fault. That would be 1/3 (e^3x) times that by e^1 which is just 1/3(e^3x)

zepdrix (zepdrix):

So we're not writing it as 2 integrals. Just rewrite the exponential using exponent rules.\[\huge \int\limits e^{3x+1}dx \quad =\quad \int\limits e^{3x}\cdot e^1dx\] See how I just rewrote the exponential term within the integral? :O

zepdrix (zepdrix):

That extra E is throwing you off I think :D hah

OpenStudy (anonymous):

Ugh probably!

zepdrix (zepdrix):

It's just a constant, so let's pull it outside of the integral.\[\huge e \int\limits e^{3x}dx\]

zepdrix (zepdrix):

Now apply the rule for integrating exponentials at this point :D Ignore the e on the outside, it will just be put back in after you integrate.

OpenStudy (anonymous):

Soo....when we integrate the 3x that would just be 1/4(x^4) right? As its' antiderivative?

OpenStudy (anonymous):

Ugh! Wait! I think I messed up! ....Crap....Wait, so if we pulled out the e constant, where did that other e come from?

zepdrix (zepdrix):

Hmmm :O

zepdrix (zepdrix):

From the Rule of Exponents :O We have a SUM within the exponent. So we can rewrite it as the PRODUCT of 2 different exponentials with the same base. Confused about this part?\[\huge e^{3x+1} \quad = \quad e^{3x}\cdot e^1\]That's where the extra E is coming from

OpenStudy (anonymous):

Ahhh...okay, so after pulling out that e we put it outside the integral. Okay...and then we just integrate the e^3x which would be....1/3 (3^3x) ....Correct?

OpenStudy (anonymous):

Crap i meant e^3x

zepdrix (zepdrix):

Yah looks good :) then we can just bring the E back inside after that. We could even put it back into the exponent if we wanted to.\[\large e\int\limits e^{3x}dx \quad = \quad e\frac{1}{3}e^{3x}\quad = \quad \frac{1}{3}e^{3x+1}\]

OpenStudy (anonymous):

Great! Thank you so much!! I think I'm getting this now. Strange how they never showed us this technique in school .

zepdrix (zepdrix):

Actually for something that, with the +1.. THAT is probably a good candidate for a U-sub. Doing these silly little exponent rules doesn't really seem worth it :D But I dunno...

OpenStudy (anonymous):

Well I'll try it later with the U sub and see if I get the same answer. That was a good practice though, thank you.

OpenStudy (anonymous):

Do you mind if I ask this one last question, I apologize if I'm taking up so much of your time.

zepdrix (zepdrix):

sure :D

OpenStudy (anonymous):

Okay so I have this problem...I think I may know how to solve this...but I Just wanna make sure I"m doing it correctly...So we have